The monopolistic competitor faces a demand curve given by q(p)=50-5p.its cost function is c(y)=4y .what is it's optimal level of output and price
1
Expert's answer
2021-07-19T03:27:01-0400

Given the demand function, the inverse demand can be written as:


"Q(p)=50\u22125p"


"Y= 50-5p"


"5p=50-Y"


"P=10-\frac{1}{5}Y"


The cost function is given as:


"c(Y) = 4Y"


Thus, given the demand and cost, the profit function can be written as:


Profit(π) = Total Revenue−Total Cost

"Profit [π]= (10−\frac{1}{5}Y)Y−4Y"


Differentiating the profit function to solve for profit-maximizing output:



"=> \frac{∂π}{∂Y}=10−\frac{2}{5}Y−4=0"


"=>6−\frac{2}{5}Y−4=0"


"=>\frac {2}{5}Y=6"


"Y=\frac{6×5}{2}=15"



Hence, the profit-maximizing output is 15 units. 


Given the output, it can be substituted in the inverse demand function to find the price:


"p=10−[\frac{1}{5}×15]"


"p =10\u22123"


"p=7"


The optimal price for the monopolist competitor to charge is 7.